The sum of the ages of two boys is 18. If 4 times the younger boy's age is subtracted from 3 times the older boy's age, the difference is 12. What are the ages of the two boys?
Answer by lwsshak3(11628)

The sum of the ages of two boys is 18. If 4 times the younger boy's age is subtracted from 3 times the older boy's age, the difference is 12. What are the ages of the two boys?
**
let x=older boy's age
18-x=younger boy's age
..
3x-4(18-x)=12
3x-72+4x=12
7x=84
x=12
18-x=6
..
older boy's age=12
younger boy's age=6